it is estimated that 30% of all drivers have medical aid. What is the probability that in a sample of 10 drivers:
A)Exactly 4 people will have a medical aid
B)At least 2 people will have a medical aid
C)More than 9 will have a medical aid
1
Expert's answer
2019-04-26T10:59:09-0400

"A)\;P(X=4)=C_{10}^4(0.3)^4(0.7)^6=0.2001\\
B)\;P(X\geq{2})=1-P(X<2)=1-P(X=0)-P(X=1)=1-C_{10}^0(0.3)^0(0.7)^{10}-C_{10}^1(0.3)^1(0.7)^9=0.8507\\
C)\;P(X>9)=P(X=10)=C_{10}^{10}(0.3)^{10}(0.7)^0=0.000006"